rsbrooks01:

a ball is thrown an initial height of 1 meter with an initial upward velocity of 5 m/s. The ball's height h (in meters) after t seconds

justjm:

use your kinematics equations to make a function between y & t. \(y_o\)=1 m \(v_{yi}\)=5m/s \(a\)=-9.8m/s^2

justjm:

You would use this one \(Δy=v_ot+\frac{1}{2}at^2\) apply it \(y=5t+\frac{1}{2}(-9.81)t^2\) If you are given a diff acceleration other than gravity then use that